OpenStudy (anonymous):

the factors of 2y2 - 2y - 4? The choices are: (2y - 4)(y - 1) (2y - 4)(y + 1) (2y - 2)(y - 2) (2y - 2)(y + 2)

OpenStudy (jadeishere):

Plug it in if it were (2y -4)(y - 1) it'd be 2y2 - 4y -2y + 4 so that'd be 2y2 - 6y + 4. So not number one

OpenStudy (anonymous):

Yeah I got that far but then I got off track!

OpenStudy (jadeishere):

(2y - 4)(y + 1) = 2y2 - 4y + 2y + 4 = 2y2 - 2y - 4 So. what do you think the answer is?

OpenStudy (anonymous):

3? I think...

OpenStudy (anonymous):

Well Im not sure if you would add or subtract in the second set of ( )

OpenStudy (jadeishere):

Okay. The 3rd and 4th one cannot be the answer because 2y* 2y = 4y. -4y + 4y = 0 so let's look at the second again?

OpenStudy (jadeishere):

with (2y - 4) (y + 1) you multiply 2y * y = 2y2

OpenStudy (jadeishere):

then multiply 2y * 1 = 2y then -4 * y = -4y -4 * 1 = -4 so add all of the together and you get 2y2 + 2y - 4y - 4

OpenStudy (anonymous):

OK thaank u!!!

OpenStudy (jadeishere):

simplify that by adding like terms and you get 2y2 = 2y - 4 :)

OpenStudy (anonymous):

! Thanks soo much!!!!

OpenStudy (jadeishere):

Yep :)